A triangle has sides with lengths in centimeters that are consecutive even integers. Find the
Length if the perimeter is 186 cm.

Answer by mark_henrhiz(40)

n
n+2
n+4


n + n+2 + n+4 = 186
3n + 6 = 186
3n = 186 -6
3n = 180
3n/3 = 180/3
n = 60


n=60
n+2=62
n+4=64
the lengths of the sides of the triangle are 60 cm, 62 cm, and 64 cm
check:
60 + 62 +64 = 186